OpenStudy (anonymous):

A company has two manufacturing plants. Plant A and Plant B. Three times Plant A's production is equal to one half of the total produced by the company. If T represents the total produced by the company, which of the following is equal to the amount produced by Plant B? A. 5/6 T B. 2/3 T C. 1/2 T D. 1/3 T E. 1/6 T I get choice A, but not sure if it's correct.

OpenStudy (anonymous):

T is the total production: So, here A's production is 1/2 of A,So we can write it as: \[A's + B's = T\] \[\frac{1}{2}T + B's = T\] Now subtract 1/2(T) from both the sides, can you do it??

OpenStudy (anonymous):

*So here A's production is 1/2 (T) sorry I have written wrong there..

OpenStudy (anonymous):

@mshi9696 reply if you can do it or not?? I am here to make you understand each and every thing but please just interact properly..

OpenStudy (anonymous):

Sorry, i was trying to answer another problem, i will be back in one minute.

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

subtract 1/T from both sides? isn't 1/2x-x=-b -1/2x=-b ....

OpenStudy (anonymous):

No.. How??

OpenStudy (anonymous):

This point you got? \[A's + B's = T\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now what is A's in terms of T??

OpenStudy (anonymous):

A's = \(\frac{1}{2}\)T right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

3a=1/2 T

OpenStudy (anonymous):

Three times Plant A's production is equal to one half of the total produced by the company.

OpenStudy (anonymous):

Oh really sorry, I misread the question. yes you are pretty right.. \(3A's = \frac{1}{2}T\)

OpenStudy (anonymous):

So find here A's by dividing it by 3..

OpenStudy (anonymous):

a=1/6

OpenStudy (anonymous):

Of T.. Don't forget..

OpenStudy (anonymous):

Oh yeah

OpenStudy (anonymous):

So: Here we go: \[A's + B's = T \implies \frac{1}{6}T + B's = T\]

OpenStudy (anonymous):

so B is 5/6 ?

OpenStudy (anonymous):

Yep... But why this is not the answer??

OpenStudy (anonymous):

No I meant that I am not sure if I was correct.

OpenStudy (anonymous):

Oh sorry you have written there that you are not sure about this, I thought you said this is ot right..

OpenStudy (anonymous):

So you are right, Have Faith in yourself..

OpenStudy (anonymous):

I don't have the answer key. :( Anyways, thank you so much!!

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (anonymous):

The way you solved I think you are the answer key yourself..